# Transversely projective structures on Smooth Foliations on Surfaces

### v1

@@ 4.1. Definition
It follows from the definition, that the reference fibration of a smooth turbulent foliation is automatically quasi-smooth. Indeed, as explained in [brunella-00-MR1948251, Section 4.3], if the support of a fiber of $f$ is not smooth then it is not irreducible and one its components is invariant by $\mathcal{F}$. Moreover, since such component have negative self-intersection, Camacho-Sad formula implies that it must contain a singularity of the foliation. But this contradicts the smoothness of $\mathcal{F}$.
@@ 4.3. Multiple and invariant fibers
The transversality of $\mathcal{F}$ and the general fiber of $f$ implies that away that the restriction of $f$ to $\pi^{-1}(B^{\circ})$ is locally trivial. Moreover, over discs $\mathbb{D}\subset B^{\circ}$ we have that $f^{-1}(\mathbb{D})\simeq\mathbb{D}\times E$ and the foliation $\mathcal{F}$ is either conjugated to the horizontal foliation defined by $dz$, when none of the fibers are $\mathcal{F}$-invariant, or, there exists a meromorphic $1$-form $\alpha$ on $\mathbb{D}$ with polar divisor equal to the union of the points below the $\mathcal{F}$-invariant fibers such that ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{f^{-1}(\mathbb{D})}}$ is defined by $\alpha+dz$.
@@
where $\tau\in E$ is a torsion point of order $n$, according to [bpv-MR2030225, Section 13 of Chapter V]. The foliation $\mathcal{F}$ is like in the smooth case except for the extra condition, in the case $f^{-1}(0)$ is a $\mathcal{F}$-invariant fiber, that the $1$-form $\alpha$ is invariant by the autmorphism $x\mapsto\exp(2\pi\sqrt{-1}/n)x$.
@@ 4.4. Transversely projective structures
Assume first that $\mathcal{F}$ is transversely projective. After passing to a finite étale we can assume that the reference fibration has no multiple fibers (Corollary 3.5) and that $\mathcal{F}$ is defined by a closed meromorphic $1$-form $\omega$ without zeros (Lemma 4.2). Let $F_{1},\ldots,F_{k}$ the $\mathcal{F}$-invariant curves that appear in $(\omega)_{\infty}$. Each $F_{i}$ corresponds to a fiber $f^{-1}(p_{i})$. Let $U\subset B$ be a simply connected open set of $B$, such that $p_{i}\in U$ for every $U$. Let $X_{U}=f^{-1}(U)$. By Ehresmann’s Theorem, $X_{U}$ is $C^{\infty}$-diffeomorphic to the product of an arbitrary fiber and $U$. In particular, $\pi_{1}(X_{U})$ is abelian. The foliation ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$ is still turbulent and transversely projective, and we are going to prove that ${\left.\kern-1.2pt\omega\vphantom{\big{|}}\right|_{X_{U}}}$ is of type (1) or (2).
Let $\mathcal{P}=(\pi:P\rightarrow X_{U},\mathcal{G},\alpha:X_{U}\rightarrow Y)$ be a projective structure for ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$, and $\rho:\pi_{1}(X_{U})\rightarrow\mathrm{PSL}(2,\mathbb{C})$ be its monodromy representation. Because $\pi_{1}(X_{U})$ is abelian, the representation $\rho$, up to conjugation, fits into one the following descriptions.
@@
$\sigma^{*}(\eta)=F\cdot{\left.\kern-1.2pt\omega\vphantom{\big{|}}\right|_{X_{U}}}$
Because $\eta$ and $\omega$ are both closed, then $F$ is first integral for ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$. Therefore, $F$ is constant by Lemma 4.3. The transversality between $\sigma$ and $\mathcal{G}$ implies that the order of poles and residues of ${\left.\kern-1.2pt\omega\vphantom{\big{|}}\right|_{X_{U}}}$ are the same as the ones of $\eta$. The result follows.
Suppose now that there exists a finite étale covering $\pi:Y\to X$ such that ~$\tilde{\mathcal{F}}=\pi^{*}\mathcal{F}$ is defined by a closed meromorphic 1-form $\omega$ of type (1) or (2). Without loss of generality, we can assume that $\pi$ is Galois with Galois group $G$. In case (1), local primitives $\int\omega$ form a collection of distinguished first integrals for $\tilde{\mathcal{F}}$ showing that $\tilde{\mathcal{F}}$ is transversely projective. Similarly, in case (2), we consider the local first integrals defined by $\exp\left(\lambda^{-1}\int\omega\right)$ to produce a transversely projective structure for $\tilde{\mathcal{F}}$.
Since $\tilde{\mathcal{F}}$ is defined by a closed meromorphic $1$-form without zeros and with non-empty polar divisor, $N_{\tilde{\mathcal{F}}}=\mathcal{O}_{Y}((\omega)_{\infty})$ is effective and non-zero. Therefore, $H^{0}((N^{*}_{\tilde{\mathcal{F}}})^{\otimes 2})=0$, and thus by Proposition 2.3, the transversely projective structure of $\tilde{\mathcal{F}}$ is unique. In particular, it is invariant by the Galois group $G$. Thus, the transversely projective structure descends to a transversely projective structure for $\mathcal{F}$. The theorem follows. ∎
@@ 5.1. Foliations on Inoue surfaces
An Inoue surface $X$ is a quotient of $\mathbb{H}\times\mathbb{C}$ by a (non-abelian) subgroup of $\operatorname{Aut}(\mathbb{H}\times\mathbb{C})\subset\operatorname{Aut}(\mathbb{C}^{2})$ generated by one diagonal linear map and three translations, see [bpv-MR2030225, Chapter V, Section 19]. The foliations defined by the natural projections to $\mathbb{H}$ and $\mathbb{C}$ descend to foliations on $X$ and are the unique foliations on $X$. Both are clearly transversely projective (actually, they are transversely affine).
@@ 5.2. Foliations on Hopf surfaces
A Hopf surface $X$ is a compact complex surface which has as universal covering $W=\mathbb{C}^{2}-\{(0,0)\}$, see [bpv-MR2030225, Chapter V, Section 18]. If $\pi_{1}(X)$ is isomorphic to $\mathbb{Z}$ then $X$ is called a primary Hopf surface, otherwise $X$ is a secondary Hopf surface. Any secondary Hopf surface admits a finite étale covering which is a primary Hopf surface. Primary Hopf surfaces in their turn, are $C^{\infty}$-diffeomorphic to $S^{3}\times S^{1}$ and biholomorphic to the cyclic quotient of $W$ by the group generated by either $\varphi(x,y)\mapsto(\alpha x,\beta y)$ where $\alpha,\beta$ are complex numbers satisfying $10<|\alpha|\leq|\beta|<$ 1 or $\varphi(x,y)\mapsto(\alpha^{n}x+\tau\cdot y^{n},\alpha y)$, where $\alpha$ is as before, $\tau\in\mathbb{C}^{*}$, and $n\geq 2$ is an integer.
@@ 5.3. Linear primary Hopf surfaces
Assume that we are in the first case, that is $\varphi:W\to W$ is a linear contraction. If the multiplicative semi-group of $\mathbb{C}^{*}$ generated by $\alpha$ and $\beta$ has rank two then the corresponding primary Hopf surface is of algebraic dimension zero and the foliations on it are quotients of the foliations defined by the $1$-forms $\lambda\frac{dx}{x}+\mu\frac{dy}{y}$ where $\lambda$ and $\mu$ are complex numbers, not both zero. If instead the multiplicative semi-group of $\mathbb{C}^{*}$ generated by $\alpha$ and $\beta$ has rank one, i.e. there exists relatively prime positive integers $n,m>0$ such that $\alpha^{n}=\beta^{m}$ then the algebraic dimension of $X$ is equal to one. Furthermore, the meromorphic function $x^{n}/y^{m}$ descends to a meromorphic function $f$ on $X$ and is such that $\mathbb{C}(X)=\mathbb{C}(f)$. Notice that $f$ has no indeterminacy points and defines a morphism from $X\to\mathbb{P}^{1}$ with elliptic fibers. Note that the orbifold structure on $\mathbb{P}^{1}$ is good only when $n=m=1$. In any case, every other foliation $\mathcal{F}$ on the primary Hopf surface $X$ is turbulent with respect to this fibration. Moreover, they are all defined by a closed meromorphic $1$-form $\omega$. We point out that $\mathcal{F}$ has at least one invariant fiber and hence, by Lemma 4.3, it does not have a meromorphic first integrals. Alternatively, this also follows from the fact the algebraic dimension of $X$ is equal to one. As a consequence, we obtain the $1$-form $\omega$ is the unique, up to multiplication by a non-zero constant, closed meromorphic $1$-form defining $\mathcal{F}$.
@@ 5.4. Non-linear primary Hopf surfaces
We now discuss the second possibility for $\varphi$, that is $\varphi$ is non-linear and of the form $\varphi(x,y)\mapsto(\alpha^{n}x+\tau\cdot y^{n},\alpha y)$. In this case, the primary Hopf surface $X$ has algebraic dimension and the foliations on $X$ are obtained as quotients of the foliations on $W$ defined by the meromorphic $1$-forms

### v2

@@ 4.1. Definition
It follows from the definition, that the reference fibration of a smooth turbulent foliation is automatically quasi-smooth. Indeed, as explained in [5, Section 4.3], if the support of a fiber of $f$ is not smooth then it is not irreducible and one of its components is invariant by $\mathcal{F}$. Moreover, since such component have negative self-intersection, Camacho-Sad formula implies that it must contain a singularity of the foliation. But this contradicts the smoothness of $\mathcal{F}$.
@@ 4.3. Multiple and invariant fibers
The transversality of $\mathcal{F}$ and the general fiber of $f$ implies that the restriction of $f$ to $\pi^{-1}(B^{\circ})$ is locally trivial. Moreover, over discs $\mathbb{D}\subset B^{\circ}$ we have that $f^{-1}(\mathbb{D})\simeq\mathbb{D}\times E$ and the foliation $\mathcal{F}$ is either conjugated to the horizontal foliation defined by $dz$, when none of the fibers are $\mathcal{F}$-invariant, or, there exists a meromorphic $1$-form $\alpha$ on $\mathbb{D}$ with polar divisor equal to the union of the points below the $\mathcal{F}$-invariant fibers such that ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{f^{-1}(\mathbb{D})}}$ is defined by $\alpha+dz$.
@@
where $\tau\in E$ is a torsion point of order $n$, according to [1, Section 13 of Chapter V]. The foliation $\mathcal{F}$ is like in the smooth case except for the extra condition, in the case $f^{-1}(0)$ is a $\mathcal{F}$-invariant fiber, that the $1$-form $\alpha$ is invariant by the autmorphism $x\mapsto\exp(2\pi\sqrt{-1}/n)x$.
@@ 4.4. Transversely projective structures
Assume first that $\mathcal{F}$ is transversely projective. After passing to a finite étale cover we can assume that the reference fibration has no multiple fibers (Corollary 3.5) and that $\mathcal{F}$ is defined by a closed meromorphic $1$-form $\omega$ without zeros (Lemma 4.2). Let $C_{1},\ldots,C_{k}$ the $\mathcal{F}$-invariant curves that appear in $(\omega)_{\infty}$. Each $C_{i}$ corresponds to a fiber $f^{-1}(p_{i})$. Let $U\subset B$ be a simply connected open set of $B$, such that $p_{i}\in U$ for every $U$. Let $X_{U}=f^{-1}(U)$. By Ehresmann’s Theorem, $X_{U}$ is $C^{\infty}$-diffeomorphic to the product of an arbitrary fiber and $U$. In particular, $\pi_{1}(X_{U})$ is abelian. The foliation ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$ is still turbulent and transversely projective, and we are going to prove that ${\left.\kern-1.2pt\omega\vphantom{\big{|}}\right|_{X_{U}}}$ is of type (1) or (2).
Let $\mathcal{P}=(\pi:Y\rightarrow X_{U},\mathcal{G},\alpha:X_{U}\rightarrow Y)$ be a projective structure for ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$, and $\rho:\pi_{1}(X_{U})\rightarrow\mathrm{PSL}(2,\mathbb{C})$ be its monodromy representation. Because $\pi_{1}(X_{U})$ is abelian, the representation $\rho$, up to conjugation, fits into one the following descriptions.
@@
$\alpha^{*}(\eta)=F\cdot{\left.\kern-1.2pt\omega\vphantom{\big{|}}\right|_{X_{U}}}.$
Because $\eta$ and $\omega$ are both closed, then $F$ is first integral for ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$. Therefore, $F$ is constant by Lemma 4.3. The transversality between $\alpha$ and $\mathcal{G}$ implies that the order of poles and residues of ${\left.\kern-1.2pt\omega\vphantom{\big{|}}\right|_{X_{U}}}$ are the same as the ones of $\eta$. The result follows.
Suppose now that there exists a finite étale covering $\pi:\tilde{X}\to X$ such that ~$\tilde{\mathcal{F}}=\pi^{*}\mathcal{F}$ is defined by a closed meromorphic 1-form $\omega$ of type (1) or (2). Without loss of generality, we can assume that $\pi$ is Galois with Galois group $G$. In case (1), local primitives $\int\omega$ form a collection of distinguished first integrals for $\tilde{\mathcal{F}}$ showing that $\tilde{\mathcal{F}}$ is transversely projective. Similarly, in case (2), we consider the local first integrals defined by $\exp\left(\lambda^{-1}\int\omega\right)$ to produce a transversely projective structure for $\tilde{\mathcal{F}}$.
Since $\tilde{\mathcal{F}}$ is defined by a closed meromorphic $1$-form without zeros and with non-empty polar divisor, $N_{\tilde{\mathcal{F}}}=\mathcal{O}_{\tilde{X}}((\omega)_{\infty})$ is effective and non-zero. Therefore, $H^{0}((N^{*}_{\tilde{\mathcal{F}}})^{\otimes 2})=0$, and thus by Proposition 2.3, the transversely projective structure of $\tilde{\mathcal{F}}$ is unique. In particular, it is invariant by the Galois group $G$. Thus, the transversely projective structure descends to a transversely projective structure for $\mathcal{F}$. The theorem follows. ∎
The next example illustrates how the theorem above, applied in some special cases, provides examples of turbulent foliations that admit (and does not admit) transversely projective structures.
Example 4.5.
Let $E$ be an elliptic curve, and let $X=E\times\mathbb{P}^{1}$ with projections $\pi_{1}:X\rightarrow E$ and $\pi_{2}:X\rightarrow\mathbb{P}^{1}$. The smooth turbulent foliations on $X$ are given by closed meromorphic 1-forms of the form $\pi_{1}^{*}(\omega_{1})+\pi_{2}^{*}(\omega_{2})$, where $\omega_{1}$ is a holomorphic non-vanishing 1-form on $E$, and $\omega_{2}$ is a meromorphic 1-form on $\mathbb{P}^{1}$ (see [11, Theorem 1.2]). Then, $\mathcal{F}$ is smooth transversely projective if and only if $\omega_{2}$ satisfies one the conditions of Theorem 4.4. For instance, considering a coordinate $z$ for $\mathbb{P}^{1}$, both $\omega_{2}=dz$ and $\omega_{2}=dz/z$ provide examples of transversely projective foliations, while $\omega_{2}=dz/z^{3}$ provides an example of a turbulent foliation without transversely projective structure. More generally, one can verify that a generic smooth turbulent foliation in $X$ does not admit a transversely projective structure.
@@ 5.1. Foliations on Inoue surfaces
An Inoue surface $X$ is a quotient of $\mathbb{H}\times\mathbb{C}$ by a (non-abelian) subgroup of $\operatorname{Aut}(\mathbb{H}\times\mathbb{C})\subset\operatorname{Aut}(\mathbb{C}^{2})$ generated by one diagonal linear map and three translations, see [1, Chapter V, Section 19]. The foliations defined by the natural projections to $\mathbb{H}$ and $\mathbb{C}$ descend to foliations on $X$ and are the unique foliations on $X$. Both are clearly transversely projective (actually, they are transversely affine).
@@ 5.2. Foliations on Hopf surfaces
A Hopf surface $X$ is a compact complex surface which has as universal covering $W=\mathbb{C}^{2}-\{(0,0)\}$, see [1, Chapter V, Section 18]. If $\pi_{1}(X)$ is isomorphic to $\mathbb{Z}$ then $X$ is called a primary Hopf surface, otherwise $X$ is a secondary Hopf surface. Any secondary Hopf surface admits a finite étale covering which is a primary Hopf surface. Primary Hopf surfaces in their turn, are $C^{\infty}$-diffeomorphic to $S^{3}\times S^{1}$ and biholomorphic to the cyclic quotient of $W$ by the group generated by either $\varphi(x,y)=(\alpha x,\beta y)$ where $\alpha,\beta$ are complex numbers satisfying $10<|\alpha|\leq|\beta|<$ 1 or $\varphi(x,y)\mapsto(\alpha^{n}x+\tau\cdot y^{n},\alpha y)$, where $\alpha$ is as before, $\tau\in\mathbb{C}^{*}$, and $n\geq 2$ is an integer.
@@ 5.3. Linear primary Hopf surfaces
Assume that we are in the first case, that is $\varphi:W\to W$ is a linear contraction. If the multiplicative semi-group of $\mathbb{C}^{*}$ generated by $\alpha$ and $\beta$ has rank two then the corresponding primary Hopf surface is of algebraic dimension zero and the foliations on it are quotients of the foliations defined by the $1$-forms $\lambda\frac{dx}{x}+\mu\frac{dy}{y}$ where $\lambda$ and $\mu$ are complex numbers, not both zero. If instead the multiplicative semi-group of $\mathbb{C}^{*}$ generated by $\alpha$ and $\beta$ has rank one, i.e. there exists relatively prime positive integers $n,m>0$ such that $\alpha^{n}=\beta^{m}$ then the algebraic dimension of $X$ is equal to one. Furthermore, the meromorphic function $x^{n}/y^{m}$ descends to a meromorphic function $f$ on $X$ and is such that $\mathbb{C}(X)=\mathbb{C}(f)$. Notice that $f$ has no indeterminacy points and defines a morphism from $X\to\mathbb{P}^{1}$ with elliptic fibers. Note that the orbifold structure on $\mathbb{P}^{1}$ is good only when $n=m=1$. In any case, every other foliation $\mathcal{F}$ on the primary Hopf surface $X$ is turbulent with respect to this fibration. Moreover, they are all defined by a closed meromorphic $1$-form $\omega$. We point out that $\mathcal{F}$ has at least one invariant fiber and hence, by Lemma 4.3, it does not have a meromorphic first integrals. Alternatively, this also follows from the fact the algebraic dimension of $X$ is equal to one. As a consequence, we obtain that the $1$-form $\omega$ is, up to multiplication by a non-zero constant, the unique closed meromorphic $1$-form defining $\mathcal{F}$.
@@ 5.4. Non-linear primary Hopf surfaces
We now discuss the second possibility for $\varphi$, that is $\varphi$ is non-linear and of the form $\varphi(x,y)\mapsto(\alpha^{n}x+\tau\cdot y^{n},\alpha y)$. In this case, the primary Hopf surface $X$ has algebraic dimension zero and the foliations on $X$ are obtained as quotients of the foliations on $W$ defined by the meromorphic $1$-forms
@@ 5.5. Synthesis
We provide an application of the preceding proposition, exhibiting cases of turbulent foliations that either possess or lack a transversely projective structure, depending on the choice of the defining 1-form.
Example 5.2.
Considering the description of the non-linear primary Hopf surfaces, Proposition 5.1 implies that a foliation $\mathcal{F}$ is transversely projective if, and only if, $\lambda=0$ and $n=1$.